Find




the Curvature and




torsion of the Curves r = (u, (1 + u)/u, (1 - u ^ 2)/u)

1
Expert's answer
2021-12-13T17:58:06-0500

The curvature of the curve given by the vector function "\bm r" is


"k(u)=\dfrac{|\bm r'(u)\times\bm r''(u)|}{|\bm r'(u)|^3}"

"\bm r'(u)=\langle 1 , \dfrac{u-(1+u)}{u^2}, \dfrac{-2u^2-(1-u^2)}{u^2}\rangle"

"=\langle 1 , -\dfrac{1}{u^2}, -(1+\dfrac{1}{u^2}\rangle"

"\bm r''(u)=\langle 0 , \dfrac{2}{u^3}, \dfrac{2}{u^3}\rangle"

"\bm r'(u)\times\bm r''(u)=\begin{vmatrix}
 \bm i & \bm j & \bm k \\
 1 & -1/u^2 & -1-1/u^2 \\
 0 & 2/u^3 & 2/u^3 \\
\end{vmatrix}"

"=(-2/u^5+2/u^3+2/u^5)\bm i-(2/u^3)\bm j+()2/u^3\bm k"

"=(2/u^3)\bm i-(2/u^3)\bm j+(2/u^3)\bm k"

"|\bm r'(u)|=\sqrt{1+1/u^4+1+2/u^2+1/u^4}"

"=\dfrac{\sqrt{2}}{u^2}\sqrt{u^4+u^2+1}"


"|\bm r'(u)\times\bm r''(u)|=\dfrac{\sqrt{6}}{u^2|u|}"

"k(u)=\dfrac{\sqrt{6}}{u^2|u|}\cdot\dfrac{u^6}{2\sqrt{2}(u^4+u^2+1)^{3/2}}"

"k(u)=\dfrac{\sqrt{3}u^2|u|}{2(u^4+u^2+1)^{3/2}}"

The torsion of the curve given by the vector function "\bm r" is


"\tau(u)=\dfrac{(\bm r'(u)\times\bm r''(u))\cdot\bm r'''(u)}{|\bm r'(u)\times\bm r''(u)|^2}"

"\bm r'''(u)=\langle 0 , -\dfrac{6}{u^4}, -\dfrac{6}{u^4}\rangle"

"(\bm r'(u)\times\bm r''(u))\cdot\bm r'''(u)=0+12/u^7-12/u^7=0"

"\tau(u)=0"